# Online conformal inference for multi-step time series forecasting

### v1

@@ 6 Conclusion and discussion
We have introduced a unified notation to formalize the mathematical representation of conformal prediction within the context of time series data, with a particular focus on multi-step time series forecasting in a generic online learning framework.
Several accessible conformal prediction methods have been extended to address the challenges of multi-step forecasting scenarios.
[…]
The methods discussed here are limited to ex-post forecasting, operating under the assumption that actual values of the exogenous predictors during the forecast period are available. In many applications, this may not be true and the methods will need to be adapted to allow forecasts of the predictors to be used, and the resulting uncertainty to be included in the conformal inference. Additionally, our methodology does not incorporate an algorithmic approach to tuning the learning rate parameter.
These considerations pave the way for numerous avenues for future research. Potential directions include the introduction of a time-dependent learning rate parameter to minimize interval width while maintaining the guaranteed coverage rate, as well as the development of refined methodologies that account for variability introduced by forecasting predictors in ex-ante scenarios.
Appendix Appendix A Proofs
A.1 Proof of Proposition 1
@@
$\hat{y}_{t+h|t}=\begin{cases}\hat{f}(y_{t},\dots,y_{t-d+1},\bm{u}_{t+1})&\text{ if }h=1,\\ \hat{f}(\hat{y}_{t+h-1|t},\dots,\hat{y}_{t+1|t},y_{t},\dots,y_{t+h-d},\bm{u}_{t+h})&\text{ if }1<h\leq d,\\ \hat{f}(\hat{y}_{t+h-1|t},\dots,\hat{y}_{t+h-d|t},\bm{u}_{t+h})&\text{ if }h>d.\\ \end{cases}$
@@
$\mbox{}\qquad+e_{t+h-1|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{1}}+\dots+e_{t+2|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{h-2}}+e_{t+1|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{h-1}}+\omega_{t+h}$
@@
$\mbox{}\qquad+e_{t+h-1|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{1}}+e_{t+h-d|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{d}}+\omega_{t+h}$
@@
[…]
A.2 Proof of Proposition 2
@@
Here, we give the proof of Proposition 2 based on Proposition 1; the idea is motivated by Sommer (2023).
@@
$\cdots$
$e_{t+h|t}$ $=\omega_{t+h}+\phi_{1}^{(h)}e_{t+h-1|t}+\cdots+\phi_{p}^{(h)}e_{t+h-p|t},\text{ with }p=\min\{d,h-1\},$
@@
[…]
A.3 Proof of Proposition 3
@@
For $T=h+1,\ldots,2h$, $E_{T}=\sum_{t=h+1}^{T}(\mathrm{err}_{t|t-h}-\alpha)\leq(T-h)-(T-h)\alpha\leq T-h\leq h\leq cg(T-h)+h$ as $c>0$, $h\geq 1$, $g$ is nonnegative, and $\mathrm{err}_{t|t-h}\leq 1$. Thus, Equation (9) holds for $T=h+1,\ldots,2h$.
@@
$\begin{cases}cg(T-h)+h-1<E_{T}\leq cg(T-h)+h,&\quad\text{ case (1) }\\ cg(T-h)+h-2<E_{T}\leq cg(T-h)+h-1,&\quad\text{ case (2) }\\ \qquad\cdots\\ cg(T-h)<E_{T}\leq cg(T-h)+1,&\quad\text{ case (h) }\\ E_{T}\leq cg(T-h).&\quad\text{ case (h+1) }\end{cases}$
In case (1), we observe that $E_{T}>cg(T-h)+h-1>cg(T-h)$, implying $q_{T+h|T}=r_{t}(E_{T})\geq b$ according to Equation (4). Thus, $s_{T+h|T}\leq q_{T+h|T}$ and $\mathrm{err}_{T+h|T}=0$. Furthermore, we have $E_{T-1}=E_{T}-(\mathrm{err}_{T|T-h}-\alpha)>cg(T-h)+h-2>cg(T-h-1)$ as $g$ is nondecreasing. This implies $q_{T+h-1|T-1}=r_{t}(E_{T-1})\geq b$, hence $s_{T+h-1|T-1}\leq q_{T+h-1|T-1}$ and $\mathrm{err}_{T+h-1|T-1}=0$. Similarly, $E_{T-2}=E_{T-1}-(\mathrm{err}_{T-1|T-h-1}-\alpha)>cg(T-h)+h-3>cg(T-h-2)$, thus $\mathrm{err}_{T+h-2|T-2}=0$. This process iterates, leading to $\mathrm{err}_{T+h|T}=\mathrm{err}_{T+h-1|T-1}=\cdots=\mathrm{err}_{T+1|T-h+1}=0$. Consequently,
@@
In case (2), we observe that $E_{T}>cg(T-h)+h-2>cg(T-h)$, thus $s_{T+h|T}\leq q_{T+h|T}$ and $\mathrm{err}_{T+h|T}=0$. Moving forward, we have $\mathrm{err}_{T+h|T}=\mathrm{err}_{T+h-1|T-1}=\cdots=\mathrm{err}_{T+2|T-h+2}=0$. Along with $\mathrm{err}_{T+1|T-h+1}\leq 1$, this means that
@@
In case (h+1), noticing $E_{T}\leq cg(T-h)$, and simply using $\mathrm{err}_{T+h-i|T-i}\leq 1$ for $i=0,\ldots,h-1$, we have
@@
Regarding the second part, $g(t-h)/(t-h)\rightarrow 0$ as $t\rightarrow\infty$ due to the sublinearity of the admissible function $g$. Hence the second part holds trivially. ∎
A.4 Proof of Proposition 4
@@
We set $q_{2h|h}=0$ without losing generality, the iteration $q_{t+h|t}=q_{t+h-1|t-1}+\eta(\mathrm{err}_{t|t-h}-\alpha)$ simplifies to $q_{t+h|t}=\eta\sum_{i=h+1}^{t}(\mathrm{err}_{i|i-h}-\alpha)$. Let $r_{t}(x)=\eta x$ and the admissible function $g(t-h)=b$, Equation (4) holds for $c=\frac{1}{\eta}$. Then Proposition 3 applies and we can easily derive the desired result. ∎
A.5 Proof of Proposition 5
@@
Let $q_{t+h|t}^{*}=q_{t+h|t}-\hat{q}_{t+h|t}$, then Equation (10) transforms into an update process $q_{t+h|t}^{*}=r_{t}\left(\sum_{i=h+1}^{t}(\mathrm{err}_{i|i-h}-\alpha)\right)$, which is an update with respect to $q_{t+h|t}^{*}$. Under this new framework, the nonconformity score becomes $s_{t+h|t}^{*}=s_{t+h|t}-\hat{q}_{t+h|t}$, with values ranging in $[-b,b]$, given the assumption that both $s_{t+h|t}$ and $\hat{q}_{t+h|t}$ fall within $[-\frac{b}{2},\frac{b}{2}]$. Thus, Proposition 3 can be directly applied to establish the long-run coverage achieved by the AcMCP method. ∎
Acknowledgments
Rob Hyndman and Xiaoqian Wang are members of the Australian Research Council Industrial Transformation Training Centre in Optimisation Technologies, Integrated Methodologies, and Applications (OPTIMA), Project ID IC200100009.

### v2

@@ 6 Conclusion and discussion
We introduced a unified notation for conformal prediction in time series, with a focus on multi-step forecasting within a general online learning framework. We extended several accessible conformal prediction methods to address the challenges of multi-step forecasting scenarios.
Under a general nonstationary autoregressive DGP, we showed that the optimal $h$-step-ahead forecast errors can be approximated by a linear combination of at most its lag (h1)(h-1) with respect to forecast horizon. Building on this foundation, we introduce AcMCP, a novel method that explicitly accounts for autocorrelations in multi-step forecast errors and achieves long-run coverage guarantees without assumptions on data distributional shifts. Across simulations and real data applications, AcMCP achieves coverage closer to the target within local windows while offering adaptive prediction intervals that respond effectively to varying conditions.
Our analysis is restricted to an ex-post forecasting setting, in which future values of exogenous predictors are assumed to be observed. In this setting, interval width reflects both aleatoric uncertainty, arising from intrinsic randomness in DGP, and epistemic uncertainty, stemming from a lack of knowledge about the underlying DGP (Sale et al. 2025). In many practical forecasting scenarios, however, predictors must themselves be forecast. In such settings, both types of uncertainty are amplified and interdependent: aleatoric uncertainty expands to include the variability induced by stochastic predictor forecasts, while epistemic uncertainty compounds through the propagation of uncertainty from the predictor models to the primary predictive model. Consequently, conformal inference procedures require adaptation to explicitly incorporate and propagate this joint uncertainty in order to preserve valid coverage and statistical efficiency. Additionally, our methodology does not incorporate an automated procedure for tuning the learning-rate parameter.
These limitations suggest several directions for future research. One avenue is to further improve interval efficiency while preserving coverage guarantees, with the goal of producing the narrowest possible prediction intervals. Our focus here is on methods that admit efficient online implementation through simple weighting or updating schemes, making them well suited for practical multi-step forecasting. More computationally intensive methods such as Bellman conformal inference (Yang, Candès & Lei 2024), which solve an optimization problem at each time step to trade off interval length and miscoverage, are not considered here. Integrating such optimization-based ideas with the proposed framework to further tighten intervals while retaining coverage guarantees is an important topic for future work. Another promising direction is the development of refined conformal methodologies for ex-ante forecasting, where uncertainty in future predictors should be explicitly incorporated into the inference procedure.
Acknowledgments
This work was supported by the Australian Research Council Industrial Transformation Training Centre in Optimisation Technologies, Integrated Methodologies, and Applications (OPTIMA) under Grant No. IC200100009; the Presidential Foundation of the Academy of Mathematics and Systems Science, Chinese Academy of Sciences, China, under Grant No. E555930101.
Disclosure statement
The authors report there are no competing interests to declare.
Supplementary materials
All proofs of the propositions and corollaries presented in the paper are provided in the Appendix.
Appendix
A1 Proof of Proposition 1
@@ Acknowledgments
[…]
@@
$\mbox{}\hskip 18.49988pt+e_{t+h-1|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{1}}+\dots+e_{t+2|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{h-2}}+e_{t+1|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{h-1}}+\omega_{t+h}$
@@
$\mbox{}\hskip 18.49988pt+e_{t+h-1|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{1}}+e_{t+h-d|t}\frac{\partial\hat{f}(\bm{a})}{\partial v_{d}}+\omega_{t+h}$
@@
Therefore, the forecast errors of optimal $h$-step-ahead forecasts follow an approximate AR($p$) process, where $p=\min\{d,h-1\}$. This implies that the optimal $h$-step-ahead forecast errors are at most serially correlated to lag (h1)(h-1). ∎
A2 Proof of Proposition 2
@@
Here, we give the proof of Proposition 2 based on Proposition 1.
@@
$\vdots$
$e_{t+h|t}$ $=\omega_{t+h}+\phi_{1}^{(h)}e_{t+h-1|t}+\dots+\phi_{p}^{(h)}e_{t+h-p|t},\text{ with }p=\min\{d,h-1\},$
@@
where $\theta_{i}$ is a complex combination of $\phi$ terms from the previous h1h-1 equations. So we conclude that the $h$-step-ahead forecast error sequence $\{e_{t+h|t}\}$ follows an approximate MA(h1)(h-1) process. ∎
A3 Proof of Proposition 3
@@
For $T=h+1,\dots,2h$, $E_{T}=\sum_{t=h+1}^{T}(\mathrm{err}_{t|t-h}-\alpha)\leq(T-h)-(T-h)\alpha\leq T-h\leq h\leq cg(T-h)+h$ as $c>0$, $h\geq 1$, $g$ is nonnegative, and $\mathrm{err}_{t|t-h}\leq 1$. Thus, Equation (9) holds for $T=h+1,\dots,2h$.
@@
[…]
In case (1), we observe that $E_{T}>cg(T-h)+h-1>cg(T-h)$, implying $q_{T+h|T}=r_{t}(E_{T})\geq b$ according to Equation (4). Thus, $s_{T+h|T}\leq q_{T+h|T}$ and $\mathrm{err}_{T+h|T}=0$. Furthermore, we have $E_{T-1}=E_{T}-(\mathrm{err}_{T|T-h}-\alpha)>cg(T-h)+h-2>cg(T-h-1)$ as $g$ is nondecreasing. This implies $q_{T+h-1|T-1}=r_{t}(E_{T-1})\geq b$, hence $s_{T+h-1|T-1}\leq q_{T+h-1|T-1}$ and $\mathrm{err}_{T+h-1|T-1}=0$. Similarly, $E_{T-2}=E_{T-1}-(\mathrm{err}_{T-1|T-h-1}-\alpha)>cg(T-h)+h-3>cg(T-h-2)$, thus $\mathrm{err}_{T+h-2|T-2}=0$. This process iterates, leading to $\mathrm{err}_{T+h|T}=\mathrm{err}_{T+h-1|T-1}=\dots=\mathrm{err}_{T+1|T-h+1}=0$. Consequently,
@@
In case (2), we observe that $E_{T}>cg(T-h)+h-2>cg(T-h)$, thus $s_{T+h|T}\leq q_{T+h|T}$ and $\mathrm{err}_{T+h|T}=0$. Moving forward, we have $\mathrm{err}_{T+h|T}=\mathrm{err}_{T+h-1|T-1}=\dots=\mathrm{err}_{T+2|T-h+2}=0$. Along with $\mathrm{err}_{T+1|T-h+1}\leq 1$, this means that
@@
In case (h+1), noticing $E_{T}\leq cg(T-h)$, and simply using $\mathrm{err}_{T+h-i|T-i}\leq 1$ for $i=0,\dots,h-1$, we have
@@
Regarding the second part, $g(t-h)/(t-h)\rightarrow 0$ as $t\rightarrow\infty$ due to the sublinearity of the admissible function $g$. Hence, the second part holds trivially. ∎
A4 Proof of Corollary 1
@@
We set $q_{2h|h}=0$ without losing generality, the iteration $q_{t+h|t}=q_{t+h-1|t-1}+\eta(\mathrm{err}_{t|t-h}-\alpha)$ simplifies to $q_{t+h|t}=\eta\sum_{i=h+1}^{t}(\mathrm{err}_{i|i-h}-\alpha)$. Let $r_{t}(x)=\eta x$ and the admissible function g(th)=bg(t-h)=b, Equation (4) holds for $c=\frac{1}{\eta}$. Then Proposition 3 applies and we can easily derive the desired result. ∎
A5 Proof of Corollary 2
@@
Let $q_{t+h|t}^{*}=q_{t+h|t}-\hat{q}_{t+h|t}$, then Equation (10) transforms into an update process $q_{t+h|t}^{*}=r_{t}\left(\sum_{i=h+1}^{t}(\mathrm{err}_{i|i-h}-\alpha)\right)$, which is an update with respect to $q_{t+h|t}^{*}$. Under this new framework, the nonconformity score becomes $s_{t+h|t}^{*}=s_{t+h|t}-\hat{q}_{t+h|t}$, with values ranging in [b,b][-b,b], given the assumption that both $s_{t+h|t}$ and $\hat{q}_{t+h|t}$ fall within $\frac{b}{2},\frac{b}{2}]$. Thus, Proposition 3 can be directly applied to establish the long-run coverage achieved by the AcMCP method. ∎